# The nature of ASASSN-24fw’s occultation: modelling the event as dimming by optically thick rings around a sub-stellar companion

### v1

@@ 5.5 Comparison with recent studies on ASASSN-24fw
After the event ended in mid-2025, it validated the predicted egress by Nair & Denisenko (2024). Soon thereafter, two groups Forés-Toribio et al. (2025) and Zakamska et al. (2025) reported their observations and conclusions. While both the studies agree on the large grain sized dusty component involved in the dimming event, and that ASASSN-24fw is an F-type star and evolved on a main-sequence, they differ on the interpretations of the event. But, Forés-Toribio et al. (2025) also find degeneracy in the age (and hence the mass of the star) like us. They also suggest the presence of a 0.25M M-dwarf secondary and that a circum-binary disc instead of a circum-planetary disc responsible for the dimming. Further, as they do not find any color change during eclipse (using their multi-band LCOGT and UKIRT observations), and find time-varying polarization during dimming, it was suggested that the occulting material is comprised of large (20m) carbonaceous or water ice grains rather than small silicate and CO2 grains.
Zakamska et al. (2025) undertook spectral study of the dimming and suggest that ASASSN-24fw is an evolved main sequence star and the dimming was caused by a truncated circum-planetary disc comprising of gas and dust of large sized grains. They provide mass of the occulter to be in between 0.5-19 MJ and the radius of the disc to be 0.7 au. This is similar to what we get from our lightcurve modeling. Moreover, their analysis using deep NaI D absorption lines which are in excess than that expected from stellar photosphere alone suggest the possibility of circum-stellar winds. Unlike their spectra neither we nor Forés-Toribio et al. (2025) found any metallic and H emission lines in our optical spectra. Interestingly, the optical spectrum of Forés-Toribio et al. (2025) taken on 9 October 2024 i.e. just after the star became dim, shows deep H absorption line, and our Keck LRIS spectrum taken on 29 October 2024 shows no H line. But our Keck HiRES spectrum taken on 25 January 2025 shows a H absorption line with a ’w’ shape or a central emission feature. Interestingly, Zakamska et al. (2025) find hydrogen emission lines in their Magellan and GHOST spectra taken on 3 February 2025 and 5 March 2025 respectively, after our Keck HiRES observations. As the only available pre-dimming spectra is from Gaia DR3 which also shows H absorption line, and that we have ruled out that ASASN-24fw is a YSO, indicates that the emission feature is unlikely to be coming from the star and maybe associated with the Hydrogen gas component in the occulting material which is re-emitting the absorbed stellar radiation into H region. The value of that we find in our HiRES spectrum is 8Å. This would have increased in the spectra by Zakamska et al. (2025) and indicates a transient nature of the H line.
@@ 5.6 Future plans
We also plan to compare the out-of-eclipse spectra with the pre-dimming ones and get a transmission spectrum to further characterize the composition of the occulter. However, at this stage, it is evident that the nature of ASASSN-24fw is degenerated with its age and the occulter is massive in size and has rings made materials in varying densities. The out-of-eclipse spectra would also shed light on the whether the H line comes from the stellar photosphere and if ASASSN-24fw is a YSO with thick disc kinematics and away from any star-forming region.
We also plan to conduct out-of-dimming polarimetric study of ASASSN-24fw and nearby ISM to understand the amount of polarization and the nature of the circum-stellar environment causing the infrared excess. Similarly, the sub-millimeter observations will be helpful in detecting the geometry of the circum-stellar environment of ASASN-24fw. The JWST observations using MIRI will help us to detect whether the composition is dominated by the silicate or the carbonaceous dust. Infact if the survival of the massive circumplanetary disc is indeed due to Bondi–Hoyle–Lyttleton accretion, we could also see signatures of ISM molecules in the NIRSpec and MIRI spectrum. Finally, observations using high-angular resolution instruments on large telescopes would help us to check the companionship of the faint source detected in our LCOGT images. All the observations, would be helpful to create a complete picture about ASASSN-24fw and give us insight about the evolution of such systems, existence of gas-giants with massive rings and the environments in which they form. Systems like ASASSN-24fw would be of great importance for the upcoming extremely large telescopes.
@@ 6 Conclusions
We have analyzed the recent dimming event around the star ASASSN-24fw. It previously exhibited stable brightness over decades of observation and then abruptly started noticeable dimming on 02 September 2024. The star had a persistent mid-infrared excess pointing towards the presence of a circum-stellar material. The estimated fractional luminosity of this material is 0.5%. But this excess did not exhibit any periodic variation in the pre-dimming NEOWISE lightcurve suggesting that the material is spread out around the star and is not located in clumps. The blackbody fit requires two components with temperatures of 1072 K and 387 K. From our lightcurve modelling, we suggest that this material is unlikely to be responsible for the recent dimming event.
@@
We modeled the lightcurve in two stages: (1) using a combination of Generalised Gaussian Distribution and Gaussian Process Regression called the Transit-Profile model and (2) assuming a point like planet with massive and inclined rings traversing the diameter of the star called the Occulting-disc model. Our first model suggests that the dimming actually started in March 2024 much before the actual alert was issued, and it progressed in stages. This indicates that the outer disc came in contact earlier than the main body much before the actual ingress. The different slopes of the model indicate that the density of the occulting material is varying and inversely proportional to the radial distance from the central body. While our model overfits the "micro-trends" due to the noise in the ASAS-SN lightcurve during the dimming, fitting a similar model to the ATLAS lightcurve suggests that the macro trends in the lightcurve representing various stages of dimming are significant. Our second model is a very simplified assumption of the occulter being a massive ring system around a gas-giant planet and passing across the stellar diameter. This model suggests that the inner diameter of the disc is 0.11 and the outer diameter of the disc is 16 or 0.17 au (within the hill sphere). Together this system weighs at least 3.42 and is having a constant transverse velocity of roughly 17 km/s. It also tells us that the ring system is has an opacity of 0.96 and is inclined to the orbital plane at an angle of 40o and azimuthally rotated about the z-axis at an angle of 50o. This orientation casts a massive elliptical shadow on the star that is responsible for the dimming event. This model does not fit the ingress very well because our assumptions that the planet is traversing the diameter of the star and the transverse velocity may actually be different for different layers or disc material. Further, our optool analysis using the post-dimming LCOGT data that we obtained from February to April 2025 indicates that the disc material is likely to be a mixture of gas and/or larger icy bodies.
@@
We modeled the pre-dimming SED of the star using and also the Bayesian inference code developed by del Burgo & Allende Prieto (2016, 2018). Using the mean value of reddening towards the star, gives us an age of roughly 8 Ma, mass around 1.6 M and [Teff, log(g), [Fe/H]] = [6250, 5.0, -0.5]. In the latter code we consider two possibilities that the star is in the pre-main sequence phase, or on the main-sequence. Whilst most of the stellar parameters derived from these models are similar to what we get from VOSA, we get two different ages of 7 Ma and 2 Ga. Since this analysis is based on the pre-dimming photometry, an accurate age and composition of the star will also be determined from future high resolution spectroscopic observations. We will also compare this spectrum with the post-dimming spectrum that we already have to compare the elemental abundances and thereby gaining in-depth understanding on the composition of the occulter. While two models suggest that the star is young, we find that its kinematics belong to the thick disc distribution. It is not close to any star-forming region nor it has arrived from one in the past 50 Ma. We also check the TESS lightcurve of ASASN-24fw. If ASASSN-24fw is a YSO or a young star, then its high magnetic field or star spot activity should be reflected in the TESS lightcurve. But the lack of any variability indicates that ASASSN-24fw is a older main-sequence star.
We obtained two post-dimming spectra from Keck LRIS and Keck HiRES. While they resemble an F-type main-sequence star, we do not see any emission lines. On comparing the optical spectra of Forés-Toribio et al. (2025) and Zakamska et al. (2025) we find that the H emission feature started to appear after January 2025. So the conclusion on the evolutionary status of the star can be only confirmed using future out-of-eclipse high-resolution spectral observations.
@@
Based on the massive ring size that we get, the companion is more likely to be a brown dwarf rather than a gas giant planet. As such massive circum-planetary discs should not survive beyond 100 Ma because the material would dissipate over time due to accretion, photo-evaporation, or dynamical clearing. Therefore, the existence of such a massive circum-planetary disc that is causing periodic dimming every 43.8 years means an extraordinary mechanism like a recent collisional replenishment, a anomalous stability, or Bondi–Hoyle–Lyttleton accretion processes are playing/might have playedrole. This challenges our understanding of circum-stellar disc lifetimes and brown dwarf environments and makes ASASN-24fw interesting system for future studies. If the age measured from future high-resolution spectra is consistent with the 2 Ga interpretation then the long-term existence of the circum-planetary scenario will be confirmed, and would require new theories about disc survival, debris disc dynamics, or unseen dynamical influences in mature systems.
@@
Events like ASASSN-24fw are unique, open to interpretation and very important to understand stellar, circum-stellar disc, and planetary evolution. Future observations are therefore required to solve the puzzle of the evolutionary status of these systems.
@@ Acknowledgements
We thank the referees for their suggestions and feedback to improve this manuscript. JPM acknowledges research support by the National Science and Technology Council of Taiwan under grant NSTC 112-2112-M-001-032-MY3. CdB acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) under grant WEAVE: EXPLORING THE COSMIC ORIGINAL SYMPHONY, FROM STARS TO GALAXY CLUSTERS and the European Regional Development Fund (ERDF) with reference PID2023-153342NB-I00/10.13039/501100011033, as well as from a Beatriz Galindo Senior Fellowship (BG22/00166) from the MICIU. The Universidad de La Laguna (ULL) and the Consejería de Economía, Conocimiento y Empleo of the Gobierno de Canarias are also gratefully acknowledged for the support provided to CdB (2024/347). GH acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement no. 695099). SZ acknowledges support from the Research Fellowship Program of the European Space Agency (ESA). SGD acknowledges generous support from the Ajax Foundation. This publication makes use of VOSA, developed under the Spanish Virtual Observatory (https://svo.cab.inta-csic.es) project funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00. VOSA has been partially updated by using funding from the European Union’s Horizon 2020 Research and Innovation Programme, under Grant Agreement No. 776403 (EXOPLANETS-A). This work has also made use of Astropy:444http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2022). We acknowledge with thanks the variable star observations from the AAVSO International Database contributed by observers worldwide and used in this research.

### v2

@@ 5.4 Comparison with recent studies on ASASSN-24fw
After the event ended in mid-2025, it validated the predicted egress by Nair and Denisenko (2024). Soon thereafter, two groups Forés-Toribio et al. (2025) and Zakamska et al. (2025) reported their observations and conclusions. While both the studies agree on the large grain sized dusty component involved in the dimming event, and that ASASSN-24fw is an F-type star and evolved on a main-sequence, they differ on the interpretations of the event. But, Forés-Toribio et al. (2025) also find degeneracy in the age (and hence the mass of the star) like us. They favour a model with the presence of a 0.25 M-dwarf secondary and that a circumbinary disc instead of a circumplanetary disc responsible for the dimming. Further, as they do not find any color change during eclipse (using their multi-band LCOGT and UKIRT observations), and find time-varying polarization during dimming, they suggest that the occulting material is comprised of large (20m) carbonaceous or water ice grains rather than small silicate and CO2 grains possible in circumbinary discs.
Zakamska et al. (2025) undertook spectral study of the dimming and suggest that ASASSN-24fw is an evolved main sequence star and the dimming was caused by a tidally truncated circumplanetary disc comprising of gas and dust of large sized grains. They also rule out shorter orbital periods using the DASCH survey and modern survey lightcurves. They provide mass of the occulter to be in between 19-76 and the radius of the disc to be 0.7 au. Moreover, their analysis using deep NaI D absorption lines which are in excess than that expected from stellar photosphere alone suggest the possibility of circumsecondary winds. Unlike their spectra neither we nor Forés-Toribio et al. (2025) found any metallic and H emission555Though we see a central emission feature, it is below the continuum level. So we refer to the shape of H-alpha line as w shaped feature. lines in our optical spectra. Interestingly, the optical spectrum of Forés-Toribio et al. (2025) taken on 9 October 2024 i.e. just after the star became dim, shows deep H absorption line, and our Keck LRIS spectrum taken on 29 October 2024 shows no H line. But our Keck HiRES spectrum taken on 25 January 2025 shows a H absorption line with a w shape or a central emission feature. Interestingly, Zakamska et al. (2025) find H emission line in their Magellan and GHOST spectra taken on 3 February 2025 and 5 March 2025 respectively, after our Keck HiRES observations. As the only available pre-dimming spectra is from Gaia DR3 which also shows H absorption line, and that we have ruled out that ASASSN-24fw is a YSO, indicates that the emission feature is unlikely to be coming from the star and maybe associated with the hydrogen gas component in the occulting material which is re-emitting the absorbed stellar radiation into H region. The value of that we find in our HiRES spectrum is 8Å. This is different if compared with the values by Zakamska et al. (2025) and therefore indicates a variable nature of the H line.
@@ 5.5 Future plans
We also plan to compare the out-of-eclipse spectra (shown in Figure (5)) with the during-dimming Keck HiRES and get a transmission spectrum to further characterize the composition of the occulter. However, at this stage, it is evident that ASASSN-24fw is an interesting object and that its nature is degenerate with its age and the occulter is massive in size and has rings made materials in varying densities. The post-dimming spectra already tells us that ASASSN24fw is not an YSO as the H line is in absorption. It also tells us that Ca H&K and Na Db lines are one of the deepest lines and the age indicator Li line is very weak. We would present our in-depth analysis on this in our upcoming papers.
We also plan to conduct out-of-dimming polarimetric study of ASASSN-24fw and nearby ISM to understand the amount of polarization and the nature of the circumstellar environment causing the infrared excess. Similarly, the sub-millimeter observations will be helpful in further constraining the circumstellar environment of ASASSN-24fw. The JWST observations using MIRI will help us to detect whether the composition of the circumstellar environment is dominated by the silicate or the carbonaceous dust. Infact if the survival of the massive circumplanetary disc is indeed due to Bondi–Hoyle–Lyttleton accretion, we could also see signatures of ISM molecules in the NIRSpec and MIRI spectrum. Finally, observations using high-angular resolution instruments on large telescopes would help us to check the companionship of the faint source detected in our LCOGT images. All the observations, would be helpful to create a complete picture about ASASSN-24fw and give us insight about the evolution of such systems, existence of gas-giants with optically thick rings and the environments in which they form. Systems like ASASSN-24fw would be of great importance for the upcoming extremely large telescopes.
@@ 6 Conclusions
We have analyzed the recent dimming event around the star ASASSN-24fw. It previously exhibited stable brightness over decades of observation and then abruptly started noticeable dimming on 02 September 2024. The star had a persistent mid-infrared excess pointing towards the presence of a circumstellar material. The estimated fractional luminosity of this material is 10%. This excess did not exhibit any periodic variation in the NEOWISE lightcurve prior to dimming suggesting that the material is spread out around the star and is not located in clumps. The blackbody fit requires two components with temperatures of 1100 K and 400 K. From our lightcurve modelling, we suggest that this material is unlikely to be responsible for the recent dimming event.
@@
We modeled the lightcurve in two stages: (1) using a combination of Generalised Gaussian Distribution and Gaussian Process Regression called the Transit-Profile model and (2) assuming a point like planet with massive and inclined rings traversing the diameter of the star called the Occulting-disc model. Our first model suggests that the dimming actually started in March 2024 much before the actual alert was issued, and it progressed in stages. This indicates that the outer disc came in contact earlier than the main body much before the actual ingress. The different slopes of the model indicate that the density of the occulting material is varying and inversely proportional to the radial distance from the central body. While our model overfits the "micro-trends" due to the noise in the ASAS-SN lightcurve during the dimming, fitting a similar model to the ATLAS lightcurve suggests that the macro trends in the lightcurve representing various stages of dimming are significant. Our second model is a very simplified assumption of the occulter being a massive and optically thick ring system around a gas-giant planet and passing across the stellar diameter. This model suggests that the inner diameter of the disc is 0.11 and the outer diameter of the disc is 16 or 0.17 au (within the hill sphere). Together this system weighs at least 3.42 and is having a constant transverse velocity of roughly 17 km/s. It also tells us that the ring system is has an opacity of 0.96 and is inclined to the orbital plane at an angle of 40o and azimuthally rotated about the orbital spin axis at an angle of 50o. This orientation casts a massive elliptical shadow on the star that is responsible for the dimming event. This model does not fit the ingress very well because our assumptions that the planet is traversing the diameter of the star and the transverse velocity may actually be different for different layers or disc material. Further, our OpTool analysis using the post-dimming LCOGT data that we obtained from February to April 2025 indicates that the disc material mainly consists of astronomical silicate dust grains around 0.4 to 0.6 m in size.
@@
We modeled the pre-dimming SED of the star using VOSA (Bayo et al., 2008) and also the Bayesian inference code developed by del Burgo and Allende Prieto (2016, 2018). Using the mean value of reddening towards the star, VOSA gives us an age of roughly 8 Ma, mass around 1.6 M and [Teff, log(g), [Fe/H]] = [6250, 5.0, -0.5]. In the latter code we consider two possibilities that the star is in the pre-main sequence phase, or on the main-sequence. Whilst most of the stellar parameters derived from these models are similar to what we get from VOSA, we get two different ages of 7 Ma and 2 Ga. Since this analysis is based on the pre-dimming photometry, an accurate age and composition of the star will also be determined from the post-dimming high resolution spectrum that we have obtained. We will also compare the during-dimming and post-dimming spectra to generate a transmission spectrum and measure and understand the composition of the occulter. While two models suggest that the star is young, we find that its kinematics belong to the thick disc distribution. It is not close to any star-forming region nor it has arrived from one in the past 50 Ma. We also check the TESS lightcurve of ASASSN-24fw. If ASASSN-24fw is a YSO or a young star, then its high magnetic field or star spot activity should be reflected in the TESS lightcurve. But the lack of any variability indicates that ASASSN-24fw is a older main-sequence star.
The post-dimming spectra from Keck LRIS and Keck HiRES resemble an F-type main-sequence star.
This line was in absorption before January 2025 and in emission later. Our spectrum taken in late January 2025 shows a central-emission w shaped reversal feature which suggests that the H line reversed from absorption to emission post January 2025. Further, in the post-dimming spectrum taken in October 2025, we do not find the H line in absorption. Therefore, it is most likely that the emission feature is associated with the occulter and not with the accreting circumstellar environment of the ASASSN-24fw.
@@
Based on the massive ring size that we get, the companion is more likely to be a brown dwarf rather than a gas giant planet. As such massive circumplanetary discs should not survive beyond 100 Ma because the material would dissipate over time due to accretion, photo-evaporation, or dynamical clearing, the existence of such a massive circumplanetary disc that is causing periodic dimming every 43.8 years means an extraordinary mechanism like a recent collisional replenishment, a anomalous disc stability, or Bondi–Hoyle–Lyttleton accretion processes are playing/might have playedrole. This challenges our understanding of circumstellar disc lifetimes and brown dwarf environments and makes ASASSN-24fw interesting system for future studies. If the age is consistent with the 2 Ga interpretation then the long-term existence of the circumplanetary scenario will be confirmed, and would require new theories about disc survival, debris disc dynamics, or unseen dynamical influences in mature systems.
@@
Events like ASASSN-24fw are unique, open to interpretation. Detailed analysis of such events are very important to understand stellar, circumstellar disc, and planetary evolution.
@@ Acknowledgements
We thank the referee(s) for their suggestions and feedback to improve this manuscript. JPM acknowledges research support by the National Science and Technology Council of Taiwan under grant NSTC 112-2112-M-001-032-MY3. CdB acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) under grant WEAVE: EXPLORING THE COSMIC ORIGINAL SYMPHONY, FROM STARS TO GALAXY CLUSTERS and the European Regional Development Fund (ERDF) with reference PID2023-153342NB-I00/10.13039/501100011033, as well as from a Beatriz Galindo Senior Fellowship (BG22/00166) from the MICIU. The Universidad de La Laguna (ULL) and the Consejería de Economía, Conocimiento y Empleo of the Gobierno de Canarias are also gratefully acknowledged for the support provided to CdB (2024/347). GH acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement no. 695099). SZ acknowledges support from the Research Fellowship Program of the European Space Agency (ESA). SGD acknowledges generous support from the Ajax Foundation. FK acknowledges support from the Spanish Ministry of Science, Innovation and Universities, under grant number PID2023-149918NB-I00. This work was also partly supported by the Spanish program Unidad de Excelencia María de Maeztu CEX2020-001058-M, financed by MCIN/AEI/10.13039/501100011033. This publication makes use of VOSA, developed under the Spanish Virtual Observatory (https://svo.cab.inta-csic.es) project funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00. VOSA has been partially updated by using funding from the European Union’s Horizon 2020 Research and Innovation Programme, under Grant Agreement No. 776403 (EXOPLANETS-A). During the post-dimming spectroscopic observations, ASASSSN-24fw was observed as a backup target during the CNTAC program CN2025B-33 (PI: W. Gieren, observer: B. Pilecki).
Software: This work made use of Astropy666http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2022), NumPy777https://numpy.org/doc/stable/index.html (Harris et al., 2020), SciPy888https://scipy.org/ (Virtanen et al., 2020) scientific libraries used for scientific computing, specutils999https://specutils.readthedocs.io/en/stable/ (Astropy-Specutils Development Team, 2019), extinction101010https://github.com/sncosmo/extinction (Barbary, 2021), and pyphot111111https://mfouesneau.github.io/pyphot/ (Fouesneau, 2025). We acknowledge with thanks the variable star observations from the AAVSO International Database121212https://www.aavso.org/ contributed by observers worldwide and used in this research.
@@
Appendix A Post-Dimming Optical Spectrum of ASASSN-24fw
We observed ASASSN-24fw on 13 October 2025 using the MIKE Spectrograph on the Magellan Telescope. Two exposures of 600 s each were taken at a resolving power of . The stacked, high-S/N spectrum is shown in Figure 5, with line identifications marked in the same manner as in Figure 3. The spectrum displays deep absorption in all prominent Balmer-series lines. The Ca ii H&K lines and the Na i D doublet also exhibit strong absorption. Most notably, H appears in deep absorption, and the Li i 6708 Å feature is negligible. The absence of detectable lithium and the lack of emission features indicate that ASASSN-24fw is not a young stellar object (YSO) but instead an older main-sequence star.